Project

General

Profile

NDOS CommissionMtng 20120710

Attending: Mat, Jan, Susan, Nick, Minerba, Leon, Gavin, Chad, Athans

Mat:
Detector status: Cooling on 54 more APDs on DiB 2

Jan:
List feb showed the following are regularly dropping out during running- 3-2 23, 2-1 21, 2-3 11
(Jan tried to look up the coating status but had trouble with what's available)
Report:

Monday to Tuesday
Run 13952 17 subruns start 2012-07-02 16:17:52, end 2012-07-03 09:10:28
FEB3-2:23 shutoff in subrun 0

Tuesday to Wednesday
Run 13953 24 subruns start 2012-07-03 14:36:14, end 2012-07-04 13:36:36
FEB1-1:38 shutoff in subrun 8
FEB2-1:21 shutoff in subrun 8
FEB2-3:11 shutoff in subrun 12
FEB3-2:23 shutoff in subrun 17
FEB1-1:35 shutoff in subrun 21

Wednesday to Thursday
Run 13954 21 subruns start 2012-07-04 14:34:14, end 2012-07-05 10:40:40
FEB3-1:26 shutoff in subrun 2
FEB2-3:9 shutoff in subrun 3
FEB2-3:44 shutoff in subrun 6

Thursday to Friday
Run 13956 18 subruns start 2012-07-05 16:09:06, end 2012-07-06 09:48:30
FEB2-1:21 shutoff in subrun 0
FEB1-1:38 shutoff in subrun 10

Friday to Saturday
Run 13960 24 subruns start 2012-07-06 15:16:23, end 2012-07-07 15:14:22
FEB2-1:21 shutoff in subrun 0
FEB3-1:26 shutoff in subrun 0
FEB3-2:23 shutoff in subrun 1

Saturday to Sunday
Run 13961 24 subruns start 2012-07-07 15:14:22, end 2012-07-08 15:12:23
FEB2-3:44 shutoff in subrun 8
FEB2-3:11 shutoff in subrun 19

Sunday to Monday
Run 13962 19 subruns start 2012-07-08 15:12:23, end 2012-07-09 09:47:03
FEB3-2:45 shutoff in subrun 7
FEB3-2:45 came back in subrun 11
FEB1-1:40 shutoff in subrun 15
FEB1-1:42 shutoff in subrun 15
FEB1-1:45 shutoff in subrun 15
FEB1-3:14 shutoff in subrun 15
FEB2-3:33 shutoff in subrun 15
FEB2-3:34 shutoff in subrun 15
FEB2-3:37 shutoff in subrun 15
FEB2-3:38 shutoff in subrun 15
FEB2-3:39 shutoff in subrun 15
FEB3-3:2 shutoff in subrun 15
FEB3-3:7 shutoff in subrun 15
FEB3-3:9 shutoff in subrun 15

Problems:
FEB3-2:23
FEB2-1:21
FEB2-3:11

13962 subrun 15 - This morning mass amounts (6:12 and 7:12 am )

Flaky behavior: 3-2 45 6 pixels masked -> Low occupancies

Leon:
APD Status - 50 APDs coated now -> Jason to begin test this week

Kurt (not attending): Working on a solution to bring APD by APD HV setting soon

Athans: Audible alarms from Chad

Martin:
FHEP Monitor -> Group Alarms coming into place
Synoptic web display formatting issue
Audible alarms??
Archiving?